OpenStudy (anonymous):

A store offers a 15% discount on all items in the store during a sale. All store employees get an additional 10% employee discount. In the equation below, C represents the cost of an item for an employee including tax, and x represents the original price of the item. C = 0.765x + 0.06(0.765x) Chris is an employee at the store. He plans to buy a coat during the sale. The price of the coat before the discount is $100. How much will Chris pay for the coat during the sale?

OpenStudy (ivettef365):

ok, you just have to substitute the 100 for the x, so you have C = 0.765(100) + 0.060(0.765(100))

OpenStudy (anonymous):

SO you have to multiply?

OpenStudy (ivettef365):

first multiply as follows: C = 76.50 + 0.060 (76.50)

OpenStudy (ivettef365):

now multiply again and you have C = 76.50 + 4.59

OpenStudy (ivettef365):

now just add

OpenStudy (anonymous):

$81.09?

OpenStudy (ivettef365):

good job